My question is if a coffee merchant has coffee beans that sell for $9 per
pound and $12 per pound. The two types are to be mixed to create 100 lb of a mixture that will sell for $11.25 per pound. How much of each type of bean should be used in the mixture?

Answer by checkley71(8403)

9X+12(100-X)=100*11.25
9X+1200-12X=1125
-3X=1125-1200
-3X=-75
3X=75
X=75/3
X=25 POUNDS OF $9 COFFEE BEANS
THAT LEAVES 100-25=75 POUINDS OF $12 BEANS.
PROOF
9*25+12*75=100*11.25
225+900=1125
1125=1125